# Position-Aware Depth Decay Decoding (): Boosting Large Language Model Inference Efficiency

### v1

@@ Abstract
Siqi Fan2, Xuezhi Fang1, Xingrun Xing3, Peng Han2, Shuo Shang2*, Yequan Wang1* 1Beijing Academy of Artificial Intelligence, Beijing, China 2University of Electronic Science and Technology of China, Chengdu, China 3Institute of Computing Automation, Chinese Academy of Sciences, Beijing, China
@@ 1 Introduction
LLMs have demonstrated impressive performance on various downstream tasks (e.g., text generation, question & answering, and sentiment analysis) using various evaluation protocols such as zero-shot, few-shot, and fine-tuning Touvron etΒ al. (2023). Notably, In-context learning ability allows LLMs to adapt to tasks using input-output examples without parameter updates Brown etΒ al. (2020); Todd etΒ al. (2024). However, their inference phases are very expensive due to the large number of parameters Liu etΒ al. (2023). LLMs employ multi-layer Transformers, focusing much of the computation on decoder blocks. For LLMs like Llama, inference complexity is per single inference, where is the word vector dimension, is the sequence length, and represents the number of decoder blocks Narayanan etΒ al. (2021). This shows that computational cost scales linearly with the number of decoder blocks. Therefore, many works explore the possibility of achieving dynamic depth in LLMs during inference Schuster etΒ al. (2022); Ma etΒ al. (2023).
Early studies in visual neural networks Β Bolukbasi etΒ al. (2017); Huang etΒ al. (2017) show that βEasyβ instance activates at shallower layers while βhardβ ones at deeper layers. Dynamic depth was then widely used in classification tasks with encoder-only LLM like BERT Li etΒ al. (2020); Liu etΒ al. (2020); Kong etΒ al. (2022), mainly through early exiting and skipping layers. The rise of generation tasks in decoder-only LLMs, prompts us to question: βCan we allocate different computational resources for generating tokens at different positions?" However, extending dynamic depth to generation tasks such as autoregressive decoding in Transformers presents challenges, as it requires computing the entire Transformer stack for each token Schuster etΒ al. (2022). This introduces two key challenges:
(i) Handling Missing State. Managing the missing hidden state and key-value cache for the current token becomes challenging if the previous token exited at a lower layer than . The common approach Corro etΒ al. (2024); Schuster etΒ al. (2022) is adding the missing state by copying. However, there are many variable factors affecting this copy operation are unclear (e.g., when to add the missing state).
(ii) Decision on Skipping Layer Number and Direction. While the concept of βeasyβ instances activate at shallower layers and βhardβ ones at deeper layers is appealing, this approach is incompatible with batching Huang etΒ al. (2017). The recent SkipDecode Corro etΒ al. (2024) addresses this by controlling the computational budget with batched exit points. However, choosing which layers to skip remains an open question. Currently, there are differing views on whether to skip head or tail layers at different timesteps. CALM Schuster etΒ al. (2022) uses a confidence measure to skip tail layers, while SkipDecode designs a batch exit function to skip head layers, but the reason behind this choice is unclear (Figure 1).
@@
We find that generation performance is sensitive to early decoding steps, as shown by Llama perplexity (PPL) decrease per token. We hypothesize that βDuring LLM generation, tokens predicted later have lower perplexity and thus require less computation." Previous work also Schuster etΒ al. (2022); Corro etΒ al. (2024) supports this, noting that early incorrect predictions affect subsequent tokens and that average loss decreases over time. These findings guide the design of .
The core of lies in decision-making strategy for each token (i.e., skip layer number and skip direction). Inspired by the decreasing behavior of Llama PPL per token, we design a power law decay function to decide how many layers to keep when generating token at time step. Here, is the total number of layers in the model, and is a hyperparameter that controls the decay rate. For the skip direction, previous works like Men etΒ al. (2024); Yang etΒ al. (2024) suggest skipping middle layers based on cosine similarity of hidden states. We further validate this by analyzing the information flow (e.g., MLP, attention activations) across each block during pretraining.
@@ 2 Related Work
Dynamic depth involves two methods: Early Exit (EE) and Skip layer. EE first appeared in CNN/DNN networks for visual tasks Bolukbasi etΒ al. (2017); Huang etΒ al. (2017); Teerapittayanon etΒ al. (2016). Subsequently, it was utilized in accelerating the inference of encoder-only architectures in BERT for classification tasks by Li etΒ al. (2020); Liu etΒ al. (2020); Li etΒ al. (2021); Kong etΒ al. (2022). Recently, Schuster etΒ al. (2022); Varshney etΒ al. (2023); Corro etΒ al. (2024) discuss confidence-based or batch exit EE for accelerate LM inference. Meanwhile, skip-layer dynamically omits the execution of middle layers (or modules) for any input token, facilitated by a gate function Wang etΒ al. (2018) or a binary router Zeng etΒ al. (2023) and layer pruning Kim etΒ al. (2024); Yang etΒ al. (2024); Song etΒ al. (2024). In addition, Dola and SLED Zhang etΒ al. (2024); Chuang etΒ al. (2024) improve model factuality from the perspective of dynamic depth, highlighting the potential of dynamic depth in various applications.
@@
Vanilla speculative decoding Leviathan etΒ al. (2023); Chen etΒ al. (2023) uses two models: a lightweight draft model for simple tokens and a powerful verification model for complex tokens. The draft model quickly predicts potential tokens, while the verification model checks their accuracy in parallel. To reduce model dependency, self-Speculative Decoding Zhang etΒ al. (2023); Elhoushi etΒ al. (2024), using a skip-layer version of the verification model as the draft model, thus avoiding additional model loading overhead. These skipped layers are searched via Bayesian optimization, which efficiently identifies the most suitable layer configurations for draft model. Compared to existing methods, our approach eliminates secondary model loading and token verification process, with key differences summarized across four dimensions in Table 1.
@@ 3.1 Preliminary: Dissecting Efficient Inference of LLMs
Mainstream LLMs (e.g., GPT, Llama) are rooted in the Transformer architecture Vaswani etΒ al. (2017), and pretrained with a full language modeling objective with a decoder-only structure, computing loss on all tokens.
@@
(i) Instance-aware dynamic depth poses challenges for batch processing and KV cache. Confidence-based methods like CALM Schuster etΒ al. (2022); Varshney etΒ al. (2023) when processing batch data, instances within the same batch have to wait until the last token in a batch exits before they can stop computing. This severely restricts the practical application of such techniques Corro etΒ al. (2024). (ii) Filling missing state has error propagation. At the sequence level, if the depth at which the previous token exits is earlier than that of subsequent tokens, filling of missing layersβs hidden state and KV cache for preceding tokens may cause error propagation. (iii) Uncertainty skip layer number and direction. CALM Schuster etΒ al. (2022) uses a confidence measure to skip tail layers, while SkipDecode designs a batch exit function to skip head layers. However, the rationale behind their choice is unclear.
@@ 3.2 Error Propagation on Filling Missing State
To handle these missing hidden states and KV cache, methods like CALM Elbayad etΒ al. (2019); Schuster etΒ al. (2022) adopt the approach of copying hidden states and recomputing KV cache, while SkipDecode Corro etΒ al. (2024) copies both hidden states and KV cache. Both methods may introduce error propagation. To this end, we first investigate the potential of layer skip in LLMs. Then we will analyze the impact of copied operation on performance, in addition to considering other factors. We use Llama2-7b Touvron etΒ al. (2023) and the BBH word sorting task Suzgun etΒ al. (2022) for these experiments.
@@
We control for the correctness of the predicted tokens to examine the potential of early exit in LLMs. That is, we first explored confidence-based oracles for decoder-only LLMs without regard to cost Schuster etΒ al. (2022). Due to the exact match evaluation metric for word sorting***A given predicted stringβs exact match score is 1 if it is the exact same as its reference string, and is 0 otherwise., when generating each token , we can compare each block βs output hidden state with the last block hidden state after passing layerβ β β The classification layer (lm_head) transforms decoder logits into a vocabulary-wide probability distribution using linear transformation and softmax, enabling word prediction by selecting the top probability option.. This allows us to identify the hidden stateβs Saturation Depth and obtain the corresponding confidence. This process can be represented by the formula: for each generated token in each block , .
CALM Schuster etΒ al. (2022) finds that in an 8-layer T5 encoder-decoder model, exits at an average of 1.53 layers per token without performance degradation. Similar observation also happens in decoder-only LLMs, for the initial phase, the first token required processing through all layers, but during the generation phase, approximately of the tokens were correctly predicted by around depth of the model (Refer to Table 2). This suggests that LLMs exhibit significant redundancy in depth, highlighting substantial potential for compute savings through layer skipping.
@@
Note that in the above Oracle experiment, missing hidden states are copied while missing KV caches are recalculated regardless of cost. When KV caches are copied, performance drops significantly (by about 20% on the word sorting task). Although CALM Schuster etΒ al. (2022) addresses this by recalculating KV states, this approach is impractical for decoder-only architectures due to high computational costs Corro etΒ al. (2024). Therefore, we examine how the copy operation affects generative tasks in Figure 2(a). We consider two variables: (i) The effect of decoder time steps (one token decoded at a time) on final performance when the same layers are copied, and (ii) the number of layers copied. For convenience, only tail layers are copied. The results are depicted in Figure 2(a), which shows that earlier perturbations lead to more significant performance degradation while copying more layers exacerbates performance decline. We attribute this outcome to perplexity (PPL) behavior per token in Figure 2(b). It provides a measure of uncertainty or βsurprise" about the prediction. For a token with probability , it is calculated as . At the start of the generation, limited information causes higher PPL. As more tokens are decoded, richer context lowers PPL for later tokens. Thus, we hypothesize that βDuring LLM generation, tokens predicted later have lower perplexity and thus require less computation.β
@@ 3.3 Core & Flex Layer
Modern LLMs build coarse-grained features in their head layers and develop more detailed, fine-grained representations in deeper layers, facilitated by multi-head attention mechanisms Vaswani etΒ al. (2017) and residual connections He etΒ al. (2016). To investigate representation changes in each decoder block, we visualize the input-output flow during training, as shown in Figure 3. Specifically, we calculate Cosine similarity and Euclidean distances for hidden state, perception, and attention activation. For instance, given output hidden states from block 1 and block 2 at training step , the cosine similarity is computed as , and the Euclidean distance is , where is the hidden size. Other metrics, such as perception (MLP) and attention activation (attn), follow similar calculation processes. Detailed data is provided in the supplementary materials.
Results show that while middle layers exhibit minimal changes in input-output flow over time, the head and tail layers remain distinct. We propose that the head and tail layers can be called core layers, with specific roles: (i) head layers handle abstract, fundamental features close to the embedding layer, and (ii) final layers align with the output near the classification layer. Middle layers, termed flex layers, are more adaptable. Dynamic depth like CALM Schuster etΒ al. (2022) skip tail layers, and SkipDecode Corro etΒ al. (2024) skip head layers. We argue that the middle flex layers should be skipped first, and the optimal starting point for skipping can be determined through parameter search in smaller models.
@@
Previous early exit methods require a stop signal for each token, such as training early exit classifiers Li etΒ al. (2020); Liu etΒ al. (2020) or using softmax response Schuster etΒ al. (2022). However, we believe these methods significantly increase computational load, especially since the softmax response projects the hidden state to a large output vocabulary size. Therefore, after identifying the core layers, we propose a decay method that does not require training or softmax response. SkipDecode Corro etΒ al. (2024) uses linear decay given a target speedup ratio. In contrast, we design a power-law decay function based on the power-law trend of PPL behavior in Figure 2(b). The decay rate is adjusted by controlling the decay coefficient , as illustrated in Figure 4(a). Given a model with layers, the skipped layers for each generated token are depicted in Figure 4(b).
@@ 4.1 Experiment Settings
We assess our method from two common text generation benchmarks from Hugging Face Open LLM Leaderboard with varying target lengths and domins: BBHΒ Suzgun etΒ al. (2022), comprising 23 challenging BIG-Bench tasksΒ Srivastava etΒ al. (2022), where previous language models fell short compared to human raters, necessitating multi-step reasoning and few-shot prompting without CoT; GSM8KΒ Cobbe etΒ al. (2021), designed for question answering on basic mathematical problems requiring multi-step reasoning. The average number of tokens in reference targets of evaluation datasets is detailed in Table 4.
@@
For performance evaluation, we report exact match score from the lm-harness evaluation frameworkΒ Gao etΒ al. (2023), this serves as the backend for Open LLM Leaderboard and is utilized by numerous open-source LLMsΒ Biderman etΒ al. (2023); Touvron etΒ al. (2023). We evaluate our approach under few-shot scenarios, using sample sizes of 3. Training set examples are added to . For in-context learning prompts, we use a default template: , concatenating random and samples from task-specific training sets.
Following CALMΒ Schuster etΒ al. (2022), our main efficiency metric is the average number of decoder layers used per output token, as it directly measures complexity reduction without conflating with implementation or infrastructure specific details. For reference, we convert it to average FLOPs reduction per output token Elbayad etΒ al. (2019); Narayanan etΒ al. (2021); Corro etΒ al. (2024). Taking into account the conditional checks and redundant parameter passing (e.g., token position) involved in , we also compared the actual speed of in real-world scenarios compared with Hugging Face implementation, reporting wall-clock time Dehghani etΒ al. (2021).
@@
For βs backbone, we choose widely recognized Llama 2 series, detailed in TableΒ 4. These models vary in terms of the number of parameters, ranging from 7 billion to 70 billion, and the number of layers, ranging from 32 layers to 80 layers. Compared with Llama 2 7/13B version, the 70B version employs Grouped Query AttentionΒ Ainslie etΒ al. (2023), enhancing its inference capabilities.
@@
We select Early Exit Schuster etΒ al. (2022) (i.e., CALM-DEC) and SkipDecode Corro etΒ al. (2024) for comparison with our . The original CALM (skip deep layers) was designed for the encoder-decoder architecture of T5 and does not support batch processing due their adaptive hidden state saturation policy. To facilitate comparison, we adapted the CALM early exit concept to suit decoder-only models and enable batch processing. SkipDecode (skip shallow layers) applies linear decay within the specified upper and lower bounds of the modelβs executable layers, given a target speedup ratio. For a fair comparison, we ensure that SkipDecode and our method have a comparable or greater average layer number, with other parameters (e.g., warm-up layer number and passing prompt inputs through all layers in the initial stage) consistent with the original paper. To ensure reproducibility, all experiments were conducted with a fixed random seed, controlling for randomness and enabling accurate comparisons.
@@
Recall that the universality decoder layers, our algorithmic improvements are fully compatible with traditional model acceleration techniques. This compatibility is noteworthy, especially considering that many earlier instance-wise dynamic depth methods Schuster etΒ al. (2022) did not support batching and KV cache. These improvements are applicable to various LLMs and can be easily transferred despite minor architectural and activation function differences.
@@
Itβs worth mentioning that the actual speedup may slightly vary during generation, given the impossibility of predicting the number of tokens the model will generate in advance Corro etΒ al. (2024). Additionally, the introduction of certain computations, such as conditional statements and array creation, also adds to the computational load. Therefore, to accurately assess the practical acceleration benefits of our approach, we integrated into the lm-harness evaluation framework and conducted end-to-end speed measurements for the entire GSM8K task within the framework, comparing it with the HF implementation. For the LLama 7B and 13B versions, we utilized 2xV100 GPUs with 32GB memory, setting the batch size to 4. As for the 70B model, we employed 8xA100 GPUs with 40GB memory, setting the batch size to 8. All three models were subjected to data parallelism during inference. The results are presented in Table 7, indicating that our method achieves significant acceleration in real-world scenarios, with this advantage becoming more pronounced as the batch size increases.

### v2

@@ Abstract
Siqi Fan1, Xuezhi Fang2, Xingrun Xing2,3, Peng Han1, Shuo Shang1*, Yequan Wang2* 1University of Electronic Science and Technology of China, Chengdu, China 2Beijing Academy of Artificial Intelligence, Beijing, China 3Institute of Computing Automation, Chinese Academy of Sciences, Beijing, China {sqfann, jedi.shang, tshwangyequan}@gmail.com
@@ 1 Introduction
LLMs have demonstrated impressive performance on various downstream tasks (e.g., text generation, question & answering, and sentiment analysis) using various evaluation protocols such as zero-shot, few-shot, and fine-tuning touvron2023llama. Notably, In-context learning ability allows LLMs to adapt to tasks using input-output examples without parameter updates brown2020language; todd2023function. However, their inference phases are very expensive due to the large number of parameters liu2023deja. LLMs employ multi-layer Transformers, focusing much of the computation on decoder blocks. For LLMs like Llama, inference complexity is per single inference, where is the word vector dimension, is the sequence length, and represents the number of decoder blocks DBLP:conf/sc/NarayananSCLPKV21. This shows that computational cost scales linearly with the number of decoder blocks. Therefore, many works explore the possibility of achieving dynamic depth in LLMs during inference schuster2022confident; ma2023llm.
Early studies in visual neural networks Β bolukbasi2017adaptive; huang2017multi show that βEasyβ instance activates at shallower layers while βhardβ ones at deeper layers. Dynamic depth was then widely used in classification tasks with encoder-only LLM like BERT li2020cascadebert; liu2020fastbert; kong2022accelerating, mainly through early exiting and skipping layers. The rise of generation tasks in decoder-only LLMs, prompts us to question: βCan we allocate different computational resources for generating tokens at different positions?" However, extending dynamic depth to generation tasks such as autoregressive decoding in Transformers presents challenges, as it requires computing the entire Transformer stack for each token schuster2022confident. This introduces two key challenges:
(i) Handling Missing State. Managing the missing hidden state and key-value cache for the current token becomes challenging if the previous token exited at a lower layer than . The common approach corro2024skipdecode; schuster2022confident is adding the missing state by copying. However, there are many variable factors affecting this copy operation are unclear (e.g., when to add the missing state).
(ii) Decision on Skipping Layer Number and Direction. While the concept of βeasyβ instances activate at shallower layers and βhardβ ones at deeper layers is appealing, this approach is incompatible with batching huang2017multi. The recent SkipDecode corro2024skipdecode addresses this by controlling the computational budget with batched exit points. However, choosing which layers to skip remains an open question. Currently, there are differing views on whether to skip head or tail layers at different timesteps. CALM schuster2022confident uses a confidence measure to skip tail layers, while SkipDecode designs a batch exit function to skip head layers, but the reason behind this choice is unclear (Figure 1).
@@
We find that generation performance is sensitive to early decoding steps, as shown by Llama perplexity (PPL) decrease per token. We hypothesize that βDuring LLM generation, tokens predicted later have lower perplexity and thus require less computation." Previous work also schuster2022confident; corro2024skipdecode supports this, noting that early incorrect predictions affect subsequent tokens and that average loss decreases over time. These findings guide the design of .
The core of lies in decision-making strategy for each token (i.e., skip layer number and skip direction). Inspired by the decreasing behavior of Llama PPL per token, we design a power law decay function to decide how many layers to keep when generating token at time step. Here, is the total number of layers in the model, and is a hyperparameter that controls the decay rate. For the skip direction, previous works like DBLP:journals/corr/abs-2403-03853; yang2024laco suggest skipping middle layers based on cosine similarity of hidden states. We further validate this by analyzing the information flow (e.g., MLP, attention activations) across each block during pretraining.
@@ 2 Related Work
Dynamic depth involves two methods: Early Exit (EE) and Skip layer. EE first appeared in CNN/DNN networks for visual tasks bolukbasi2017adaptive; huang2017multi; teerapittayanon2016branchynet. Subsequently, it was utilized in accelerating the inference of encoder-only architectures in BERT for classification tasks by li2020cascadebert; liu2020fastbert; li2021accelerating; kong2022accelerating. Recently, schuster2022confident; varshney2023accelerating; corro2024skipdecode discuss confidence-based or batch exit EE for accelerate LM inference. Meanwhile, skip-layer dynamically omits the execution of middle layers (or modules) for any input token, facilitated by a gate function wang2018skipnet or a binary router zeng2023learning and layer pruning kim2024shortened; yang2024laco; song2024sleb. In addition, Dola and SLED zhang2024sled; DBLP:conf/iclr/ChuangXLKGH24 improve model factuality from the perspective of dynamic depth, highlighting the potential of dynamic depth in various applications.
@@
Vanilla speculative decoding DBLP:conf/icml/LeviathanKM23; DBLP:journals/corr/abs-2302-01318 uses two models: a lightweight draft model for simple tokens and a powerful verification model for complex tokens. The draft model quickly predicts potential tokens, while the verification model checks their accuracy in parallel. To reduce model dependency, self-Speculative Decoding zhang2023draft; elhoushi2024layer, using a skip-layer version of the verification model as the draft model, thus avoiding additional model loading overhead. These skipped layers are searched via Bayesian optimization, which efficiently identifies the most suitable layer configurations for draft model. Compared to existing methods, our approach eliminates secondary model loading and token verification process, with key differences summarized across four dimensions in Table 1.
@@ 3.1 Preliminary: Dissecting Efficient Inference of LLMs
Mainstream LLMs (e.g., GPT, Llama) are rooted in the Transformer architecture DBLP:conf/nips/VaswaniSPUJGKP17, and pretrained with a full language modeling objective with a decoder-only structure, computing loss on all tokens.
@@
(i) Instance-aware dynamic depth poses challenges for batch processing and KV cache. Confidence-based methods like CALM schuster2022confident; varshney2023accelerating when processing batch data, instances within the same batch have to wait until the last token in a batch exits before they can stop computing. This severely restricts the practical application of such techniques corro2024skipdecode. (ii) Filling missing state has error propagation. At the sequence level, if the depth at which the previous token exits is earlier than that of subsequent tokens, filling of missing layersβs hidden state and KV cache for preceding tokens may cause error propagation. (iii) Uncertainty skip layer number and direction. CALM schuster2022confident uses a confidence measure to skip tail layers, while SkipDecode designs a batch exit function to skip head layers. However, the rationale behind their choice is unclear.
@@ 3.2 Error Propagation on Filling Missing State
To handle these missing hidden states and KV cache, methods like CALM elbayad2019depth; schuster2022confident adopt the approach of copying hidden states and recomputing KV cache, while SkipDecode corro2024skipdecode copies both hidden states and KV cache. Both methods may introduce error propagation. To this end, we first investigate the potential of layer skip in LLMs. Then we will analyze the impact of copied operation on performance, in addition to considering other factors. We use Llama2-7b touvron2023llama and the BBH word sorting task suzgun2022challenging for these experiments.
@@
We control for the correctness of the predicted tokens to examine the potential of early exit in LLMs. That is, we first explored confidence-based oracles for decoder-only LLMs without regard to cost schuster2022confident. Due to the exact match evaluation metric for word sorting111A given predicted stringβs exact match score is 1 if it is the exact same as its reference string, and is 0 otherwise., when generating each token , we can compare each block βs output hidden state with the last block hidden state after passing layer222The classification layer (lm_head) transforms decoder logits into a vocabulary-wide probability distribution using linear transformation and softmax, enabling word prediction by selecting the top probability option.. This allows us to identify the hidden stateβs Saturation Depth and obtain the corresponding confidence. This process can be represented by the formula: for each generated token in each block , .
CALM schuster2022confident finds that in an 8-layer T5 encoder-decoder model, exits at an average of 1.53 layers per token without performance degradation. Similar observation also happens in decoder-only LLMs, for the initial phase, the first token required processing through all layers, but during the generation phase, approximately of the tokens were correctly predicted by around depth of the model (Refer to Table 2). This suggests that LLMs exhibit significant redundancy in depth, highlighting substantial potential for compute savings through layer skipping.
@@
Note that in the above Oracle experiment, missing hidden states are copied while missing KV caches are recalculated regardless of cost. When KV caches are copied, performance drops significantly (by about 20% on the word sorting task). Although CALM schuster2022confident addresses this by recalculating KV states, this approach is impractical for decoder-only architectures due to high computational costs corro2024skipdecode. Therefore, we examine how the copy operation affects generative tasks in Figure 2(a). We consider two variables: (i) The effect of decoder time steps (one token decoded at a time) on final performance when the same layers are copied, and (ii) the number of layers copied. For convenience, only tail layers are copied. The results are depicted in Figure 2(a), which shows that earlier perturbations lead to more significant performance degradation while copying more layers exacerbates performance decline. We attribute this outcome to perplexity (PPL) behavior per token in Figure 2(b). It provides a measure of uncertainty or βsurprise" about the prediction. For a token with probability , it is calculated as . At the start of the generation, limited information causes higher PPL. As more tokens are decoded, richer context lowers PPL for later tokens. Thus, we hypothesize that βDuring LLM generation, tokens predicted later have lower perplexity and thus require less computation.β
@@ 3.3 Core & Flex Layer
Modern LLMs build coarse-grained features in their head layers and develop more detailed, fine-grained representations in deeper layers, facilitated by multi-head attention mechanisms DBLP:conf/nips/VaswaniSPUJGKP17 and residual connections he2016deep. To investigate representation changes in each decoder block, we visualize the input-output flow during training, as shown in Figure 3. Specifically, we calculate Cosine similarity and Euclidean distances for hidden state, perception, and attention activation. For instance, given output hidden states from block 1 and block 2 at training step , the cosine similarity is computed as , and the Euclidean distance is , where is the hidden size. Other metrics, such as perception (MLP) and attention activation (attn), follow similar calculation processes. Detailed data is provided in the supplementary materials.
Results show that while middle layers exhibit minimal changes in input-output flow over time, the head and tail layers remain distinct. We propose that the head and tail layers can be called core layers, with specific roles: (i) head layers handle abstract, fundamental features close to the embedding layer, and (ii) final layers align with the output near the classification layer. Middle layers, termed flex layers, are more adaptable. Dynamic depth like CALM schuster2022confident skip tail layers, and SkipDecode corro2024skipdecode skip head layers. We argue that the middle flex layers should be skipped first, and the optimal starting point for skipping can be determined through parameter search in smaller models.
@@
Previous early exit methods require a stop signal for each token, such as training early exit classifiers li2020cascadebert; liu2020fastbert or using softmax response schuster2022confident. However, we believe these methods significantly increase computational load, especially since the softmax response projects the hidden state to a large output vocabulary size. Therefore, after identifying the core layers, we propose a decay method that does not require training or softmax response. SkipDecode corro2024skipdecode uses linear decay given a target speedup ratio. In contrast, we design a power-law decay function based on the power-law trend of PPL behavior in Figure 2(b). The decay rate is adjusted by controlling the decay coefficient , as illustrated in Figure 4(a). Given a model with layers, the skipped layers for each generated token are depicted in Figure 4(b).
@@ 4.1 Experiment Settings
We assess our method from two common text generation benchmarks from Hugging Face Open LLM Leaderboard with varying target lengths and domins: BBHΒ suzgun2022challenging, comprising 23 challenging BIG-Bench tasksΒ srivastava2022beyond, where previous language models fell short compared to human raters, necessitating multi-step reasoning and few-shot prompting without CoT; GSM8KΒ cobbe2021gsm8k, designed for question answering on basic mathematical problems requiring multi-step reasoning. The average number of tokens in reference targets of evaluation datasets is detailed in Table 4.
@@
For performance evaluation, we report exact match score from the lm-harness evaluation frameworkΒ eval-harness, this serves as the backend for Open LLM Leaderboard and is utilized by numerous open-source LLMsΒ biderman2023pythia; touvron2023llama. We evaluate our approach under few-shot scenarios, using sample sizes of 3. Training set examples are added to . For in-context learning prompts, we use a default template: , concatenating random and samples from task-specific training sets.
Following CALMΒ schuster2022confident, our main efficiency metric is the average number of decoder layers used per output token, as it directly measures complexity reduction without conflating with implementation or infrastructure specific details. For reference, we convert it to average FLOPs reduction per output token elbayad2019depth; DBLP:conf/sc/NarayananSCLPKV21; corro2024skipdecode. Taking into account the conditional checks and redundant parameter passing (e.g., token position) involved in , we also compared the actual speed of in real-world scenarios compared with Hugging Face implementation, reporting wall-clock time dehghani2021efficiency.
@@
For βs backbone, we choose widely recognized Llama 2 series, detailed in TableΒ 4. These models vary in terms of the number of parameters, ranging from 7 billion to 70 billion, and the number of layers, ranging from 32 layers to 80 layers. Compared with Llama 2 7/13B version, the 70B version employs Grouped Query AttentionΒ ainslie2023gqa, enhancing its inference capabilities.
@@
We select Early Exit schuster2022confident (i.e., CALM-DEC) and SkipDecode corro2024skipdecode for comparison with our . The original CALM (skip deep layers) was designed for the encoder-decoder architecture of T5 and does not support batch processing due their adaptive hidden state saturation policy. To facilitate comparison, we adapted the CALM early exit concept to suit decoder-only models and enable batch processing. SkipDecode (skip shallow layers) applies linear decay within the specified upper and lower bounds of the modelβs executable layers, given a target speedup ratio. For a fair comparison, we ensure that SkipDecode and our method have a comparable or greater average layer number, with other parameters (e.g., warm-up layer number and passing prompt inputs through all layers in the initial stage) consistent with the original paper. To ensure reproducibility, all experiments were conducted with a fixed random seed, controlling for randomness and enabling accurate comparisons.
@@
Recall that the universality decoder layers, our algorithmic improvements are fully compatible with traditional model acceleration techniques. This compatibility is noteworthy, especially considering that many earlier instance-wise dynamic depth methods schuster2022confident did not support batching and KV cache. These improvements are applicable to various LLMs and can be easily transferred despite minor architectural and activation function differences.
@@
Itβs worth mentioning that the actual speedup may slightly vary during generation, given the impossibility of predicting the number of tokens the model will generate in advance corro2024skipdecode. Additionally, the introduction of certain computations, such as conditional statements and array creation, also adds to the computational load. Therefore, to accurately assess the practical acceleration benefits of our approach, we integrated into the lm-harness evaluation framework and conducted end-to-end speed measurements for the entire GSM8K task within the framework, comparing it with the HF implementation. For the LLama 7B and 13B versions, we utilized 2xV100 GPUs with 32GB memory, setting the batch size to 4. As for the 70B model, we employed 8xA100 GPUs with 40GB memory, setting the batch size to 8. All three models were subjected to data parallelism during inference. The results are presented in Table 7, indicating that our method achieves significant acceleration in real-world scenarios, with this advantage becoming more pronounced as the batch size increases.